HUBBLE SPACE TELESCOPE DAILY REPORT #2409 PERIOD COVERED: 0000Z (UTC) 07/07/99 - 0000Z (UTC) 07/08/99 Daily Status Report as of 189/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Two Sets of WF/PC-2/STIS/CCD 8267 (Taking the Measure of Planets in the Globular Cluster 47 Tucanae) Both the WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of the globular cluster NGC-104 (a.k.a., Globular Cluster 47 Tucanae) to search for stars with planetary companions in close orbits. HST is uniquely capable of providing the requisite, high-dynamic range, time-series photometry. These observations extended over a several day period. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 17 Successful : 17 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: The flight operations team moved to the SAMS back-up facility at 189/0531Z in preparation for GSFC Internal Simulation #7 (EVA-1) which will be conducted today. They will be using the PRS ground system. The multi-day Tucanae proposal, described in 1.2, continues with no reported problems. Using ROP SR-1, the SSR EDAC error counter was cleared three times. The STIS MCE-2 reset at 189/0227Z while the low voltage was on and while in an SAA interval. Recovery of the MAMA-2 will be accomplished via SMS commanding at 193/2357Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH